OpenStudy (anonymous):

Find the exact area below the curve y=x^(3)(1 - x) and above the x-axis?

OpenStudy (anonymous):

Integrating you get (x^4)/4-(x^5)/5 +c for the indefinite integral, and from a to b is just do the same substitution I did on the other exercise

OpenStudy (anonymous):

y(x)=x^3(1 - x) y(x)=x^3 - x^4 First, we have to find the two points at which the function intersects the x-axis. Let's evaluate the original equation to find where it is equal to zero. y = 0 x^3(1-x)=0 x=0 and x=1 At (0,0) and (1,0) Then Integrate the function, and evaluate the integral between those 2 points. \[\int\limits_{0}^{1} y(x) dx = \int\limits_{0}^{1} (x^3-x^4) dx\] = x^4/4-x^5/5 Area above x-axis = y(1)-y(0) = (1/4-1/5)-(0-0)=1/20